an ideal spring has an extension a when the tension in it is 4N.it length is b when tension is 5N.what is length when tension is 9 n

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear jauneet

let the natural length of spring is l

then F =kx

   4 = k*a

 k=4/a  

 

and agian  F=kx

             5 = k(b-l)

            5=4/a (b-l)

          b-l = 5a/4

          l= b-5a/4

 

let the length is x when tension is 9N

 9 = 4/a (x-l)

9a/4 = x- (b-5a/4)

 x = b-5a/4 +9a/4

   =b-a

 

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin